OpenStudy (anonymous):

A cars traveling initially at 7.94m/s acceleration rate of 0.953m/s^2 for 4.74s what is its velocity final

OpenStudy (anonymous):



OpenStudy (anonymous):

In order to solve this, you can use the equation Vf=Vi+at, where Vf is final velocity and Vi is initial velocity, a is acceleration, and t is time. Since you're solving for final velocity, you can just put your numbers into the equation. Vf=7.94+(.953*4.74) Vf=13.633 Hope that helps!